Tohban Report 2014-01-15

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 08 Jan 2014
End Date: 15 Jan 2014
Tohban: Hazel Bain
Tohban email:
Next Tohban: Nicole Duncan
List all reports


Solar Activity

Solar activity was low to moderate this week with 1 M class flare and 26 C class flares, the majority of which we produced by AR 11944 which produced the M and X class flares last week. This AR is now on the limb and will rotate off the disk within the next 24 hours. There are currently 7 other ARs on disk which have produced a few low C class flares over the last week. STEREO B shows several other active regions which will rotate onto the disk in the coming week

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    26     1     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    19     1     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     9     1     0

There were RHESSI flares/GOES flares 318 / 27 over the time range 08-Jan-14 15-Jan-14

Memory Management

The SSR fill level reached a maximum of 28%. The SSR is now emptying at the end of each daily pass set.

Spacecraft Status

The cold plate 1 and 2 are at about 118.9 K and 117.4 K, respectively. Cryocooler is nominal.

Data Gaps

One short data gap:

GAP START TIME GAP END TIME GAP (SEC) 2014-01-10T19:15:00.000 -- 2014-01-10T19:20:00.000 300.00000

Detector issues

Detector 9 showed an marked increase in the rear fast LLD over several days at the start of the week. In response Detector 9's rear fast threshold was raised from 0x60 to 0x70 at 011-07:01:37 UTC.

Other notes

Spacecraft Management

Decimation Normal/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Enabled, during entire eclipse
Night time data (rears) Enabled, during entire eclipse
Require extra passes? No?
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Detector 9 rear fast LLD threshold raised from 0x60 to 0x70.
Personal tools